OpenStudy (anonymous):

(1/625) power of1/4 raise to the power or take the root. assume all variables are positive???????

OpenStudy (anonymous):

\[ \left ( 1 \over 625\right )^{1 \over 4} = \left ( 1 \over 5\right )^{ 4 \times {1 \over 4}} = {1 \over 5} \]

OpenStudy (anonymous):

thank you